OpenStudy (anonymous):

Solve 9x − 5 = 5x + −37 for x.

OpenStudy (anonymous):

start by moving all the numbers with x to the left side, and all the numbers without x to the right side: so, first of all we subtract 5x from both sides: 9x - 5 - 5x = -37 so now the 5x is gone from the right side, now we add 5 to each side, to get rid of the -5 on the left side: 9x - 5x = -37 + 5 then we solve what we can already: 4x = -32 and then we divide each side by 4: x = -8 hope that helped :)